# Spaces of subgroups of toral groups

### v1

@@ 7.C.1. $W$ of order 2
Starting with $W$ of order 2, the indecomposable representations of the group of order 2 are $\mathbb{Z}$ (trivial), $\tilde{\mathbb{Z}}$ (generator acts as $-1$) and the two dimensional representation $\mathbb{Z}W$ [3, 74.3]. We have $H^{3}(W;\mathbb{Z})=H^{3}(W;\mathbb{Z}W)=0$, so extensions in those cases are split and $G=T^{2}\rtimes W$. On the other hand, $H^{3}(W;\tilde{\mathbb{Z}})=\mathbb{Z}/2$, where the non-split extension is $Pin(2)$ and $G\cong Pin(2)\times T$.
@@ 7.C.2. $W=C_{2}\times C_{2}$
Case 1: $\Lambda_{0}$ is the $A_{1}\times A_{1}$ lattice, so that $\Lambda_{0}=\mathbb{Z}^{2}=\mathrm{Cub}$ is the standard lattice where the coordinates are $(x,y)$, and $W$ acts by reflection in the axes. We therefore find $\Lambda_{0}=\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$ where ~$\tilde{\mathbb{Z}}_{x}$ is the $x$-axis and ~$\tilde{\mathbb{Z}}_{y}$ is the $y$-axis. Factoring out the reflection in the $x$-axis or the $y$-axis gives two Serre spectral sequences. Comparing the two we see the differentials are zero in degrees affecting $H^{i}(W;\mathbb{Z}_{x})$ with $i\leq 3$ and hence $H^{2}(W;\tilde{\mathbb{Z}}_{x})\cong\mathbb{Z}/2$ and $H^{3}(W;\tilde{\mathbb{Z}}_{x})\cong(\mathbb{Z}/2)^{2}$. This gives the tabulated value. We obtain examples which are products of $SU(2)$ and $SO(3)$; the large number arise since there is a choice about which element of $W$ lifts to something whose square is non-trivial in each factor.
Case 2: $\Lambda_{0}$ is the $A_{1}\times A_{1}$ lattice, so that $\Lambda_{0}=\mathbb{Z}^{2}=\mathrm{Cub}$ is the standard lattice where the coordinates are $(x,y)$, but now $W$ acts by reflection in the diagonal lines $x=\pm y$. Picking $C$ to be the subgroup generated by reflection in $x=y$, we therefore find $\Lambda_{0}|_{C}=\mathbb{Z}C$. This shows $H^{i}(C;\Lambda_{0})=H^{0}(C;\Lambda_{0})=\tilde{\mathbb{Z}}$, where the tilde indicates the action of reflection in $x=-y$. Accordingly $H^{i}(W;\Lambda_{0})\cong H^{i}(C^{\prime};\tilde{\mathbb{Z}})$ giving $H^{2}(W;\Lambda_{0})=0$ and $H^{3}(W;\Lambda_{0})\cong\mathbb{Z}/2$. This gives the tabulated value.
@@ 7.C.3. $W=C_{4}$
Here we restrict attention to the case when $\Lambda_{0}$ is the $B_{2}$ lattice, so that $\Lambda_{0}=\mathbb{Z}^{2}=\mathrm{Cub}$ is the standard lattice where the coordinates are $(x,y)$, and $W$ acts by a quarter turn rotation. We therefore find $\Lambda_{0}=\mathrm{ind}_{C_{2}}^{C_{4}}\tilde{\mathbb{Z}}$, so that $H^{i}(W;\Lambda_{0})=H^{i}(C_{2};\tilde{\mathbb{Z}})$ and $H^{2}(W;\Lambda_{0})=0$ and $H^{3}(W;\Lambda_{0})\cong\mathbb{Z}/2$. This gives the tabulated value.
@@ 7.C.4. $W=D_{8}$
Considering the extension $1\longrightarrow\langle a,b\rangle\longrightarrow D_{8}\longrightarrow C_{2}\longrightarrow 1$ we see that $H^{i}(D_{8};\Lambda_{0})$ is $022X$ where $X$ is either of order 2 or 4.
Considering the extension $1\longrightarrow Z\longrightarrow D_{8}\longrightarrow C_{2}\times C_{2}\longrightarrow 1$, there is a potential differential from $E_{2}^{0,2}$. If it were to be zero we would have $H^{2}(D_{8};\Lambda_{0})$ of order 4, which we have seen to be wrong. So the differential is zero and by the module structure the entire 2-line is the source of a $d_{2}$. Therefore $E_{2}^{0,3}$ survives, and $H^{i}(D_{8};\Lambda_{0})$ is $0222$ giving the tabulated values.
@@ 7.C.5. $W$ of order 3
$0\longrightarrow\mathbb{Z}\longrightarrow\mathbb{Z}W\longrightarrow\mathbb{Z}[\omega]\longrightarrow 0$
@@ 7.C.6. $W=D_{6}$
The one dimensional representations are again $\mathbb{Z}$ or $\tilde{\mathbb{Z}}$. If $\Lambda$ is of rank 2 then either $C_{3}$ acts trivially (and $\Lambda$ is inflated from the group $\Sigma_{3}/C_{3}$ of order 2) or else $M$ is a copy of $\mathbb{Z}[\omega]$ as a $C_{3}$-module. The reflection $\tau$ in the $x$-axis cannot act as a scalar (since $\tau g\tau=g^{-1}$), so it must have eigenvalues $+1$ and $-1$. The first example is when $\tau$ acts as complex conjugation on $\mathbb{Z}[\omega]$.
@@ 7.D. $W$ of prime order
We insert one more generic answer to hint at the complexities in higher rank. By [3, 74.3], the indecomposable representations of a group of prime order $p$ are $\mathbb{Z}$, $\mathfrak{a}$ for a fractional ideal $\mathfrak{a}$ in $\mathbb{Q}W$ (of rank $p-1$) or $\Lambda_{\mathfrak{a}}$ of rank $p$, where
@@
if $\Lambda_{\mathfrak{a}}=\mathfrak{a}\oplus\mathbb{Z}\cdot y$ additively, we may suppose $gy=1+y$. This means that $\mathbb{T}$ is a product of tori of dimension $11,p-$ 1 and $p$.
@@ 8. The case $W=C_{2}$, $\Lambda_{0}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$
We describe here the space of full subgroups as in the title. Examples of such groups $G$ include $T\times O(2)$ and $T\times Pin(2)$.
@@ 8.B. Lattices and Weyl groups
(a) The lattice $\Lambda_{+}(m)$ with basis $\lambda=(m,0)$ for an integer $m\geq 1$. As a $W$-module this is $\mathbb{Z}$. It corresponds to $S=C_{m}\times\tilde{T}$ and has Weyl group $T$.
(b) The lattice $\Lambda_{-}(n)$ with basis $\lambda=(0,n)$ for an integer $n\geq 1$. As a $W$-module this is $\tilde{\mathbb{Z}}$. It corresponds to $S=Z\times C_{n}$ and has Weyl group $1\times C_{2}$.
Proof : The only comment to make is that if $\Lambda$ is an invariant rank 1 lattice then $g\Lambda=\Lambda$ so that a generator $\lambda$ is taken to ±$\pm\lambda$. For the Weyl groups we note $M(m,0)^{t}=(0,0)^{t}$ but $M(0,n)^{t}=(0,-2n)^{t}$. ∎
@@
Proof : Every rank 2 lattice $\Lambda$ contains an element on the $x$-axis, and we may choose the smallest positive element $(m,0)$, and similarly $(0,n)$ is the smallest positive point on the $y$-axis. If $\Lambda$ is generated by $(m,0)$ and $(0,n)$ we have a Type 1 lattice. Otherwise there is another element $(a,b)$ with $0<a<m,0<b<n$. Since $g(a,b)=(a,-b)$ we find $(2a,0)$ is in the lattice so $2a=m$. If $2b$ is not a multiple of $n$ this contradicts the choice of $n$. Hence $(a,b)=(m,n)/2$ and we have a Type 2 lattice. ∎
@@ 8.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@
In the first case (corresponding to $\Lambda_{-}(n)$), $T^{2}/S$ is a circle with nontrival action so the cohomology of $T^{2}/S$ is in even degrees, and the lift of $\epsilon(G)$ is unique is unique up to conjugacy if it exists. In the split case $\epsilon(G)=0$ so the lift exists. In the non-split case, it lifts if and only if $y$ is even.
@@ 8.D. Summary
(this is the reason for indexing the lattice $\Lambda_{2}(2m,2n)$ using $(2m,2n)$, rather than $(m,n)$). Recalling $\Lambda_{2}(2m,2n)$ is generated by $(m,n),(m,-n)$ one may check that the inclusions are both cofree and that the first is cofree and $W$-cotrivial. They are therefore represented vertically and horizontally in the picture.
@@
[…]
@@
[…]
@@ 9.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@ 10.A. Preamble
We have $W$ of order 2 and the toral lattice is $\Lambda_{0}=\mathbb{Z}W$. In this case $H^{2}(W;T^{2})=H^{3}(W;\Lambda_{0})=0$, and all such extensions split. To be concrete $W$ acts on $T^{2}$ by exchanging factors: this example is realised by the normalizer of the maximal torus in $U(2)$.
@@ 10.B. Lattices and Weyl groups
(a) The lattice $\Lambda_{+}(m)$ with basis $\lambda=(m,m)=m(1+w)$ for an integer $m\geq 1$. As a $W$-module, this is $\mathbb{Z}$. It corresponds to $S=C_{m}\times T$ with $C_{m}$ central and has Weyl group $T$.
(b) The lattice $\Lambda_{-}(m)$ with basis $\lambda=(m,-m)=m(1-w)$ for an integer $y\geq 1$. As a $W$-module, this is $\tilde{\mathbb{Z}}$. This corresponds to $S=T\times C_{m}$ with $T$ being the centre and has Weyl group $1\times C_{2}$.
@@
Type 2: If $m+n$ and $m-n$ are even we have $\Lambda_{2}(m,n)$ which contains $\Lambda_{1}(m,n)$ as a sublattice of index 2. It has basis $\lambda_{1}=((m+n)/2,(m-n)/2)=((m+n)+(m-n)w)/2$ and $\lambda_{2}=w\lambda_{1}=((m-n)/2,(m+n)/2)=((m-n)+(m+n)w)/2$ and is $\mathbb{Z}W$ as a $W$-module. We have containments
@@
Proof : The eigenspaces of $\mathbb{Z}W$ are generated by $(1+w)$ and $(1-w)$. If a lattice $\Lambda$ is generated by its intersections with $\mathbb{Z}W_{+}$ and $\mathbb{Z}W_{-}$, then it is of Type 1. Otherwise there will be an additional element, and we argue as in Lemma 8.2 to see that the lattice must be of Type 2. ∎

### v2

@@ 7.C.1. $W$ of order 2
Starting with $W$ of order 2, the indecomposable representations of the group of order 2 are $\mathbb{Z}$ (trivial), $\tilde{\mathbb{Z}}$ (generator acts as 1-1) and the two dimensional representation $\mathbb{Z}W$ [3, 74.3]. We have $H^{3}(W;\mathbb{Z})=H^{3}(W;\mathbb{Z}W)=0$, so extensions if $\Lambda_{0}=\mathbb{Z}^{2}$ or $\mathbb{Z}W$ are split and $G=T^{2}\times W$ in the first case and $T^{2}\rtimes W$ in the second (with $W$ exchanging the two factors). On the other hand, $H^{3}(W;\tilde{\mathbb{Z}})=\mathbb{Z}/2$, where the non-split extension is $Pin(2)$. If $\Lambda_{0}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ this gives $G\cong O(2)\times T$ in the split case and $G=Pin(2)\times T$ in the non-split case. If $\Lambda_{0}=\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$ there are three non-split cases depending on which element of $T^{2}[2]$is a square.
@@ 7.C.2. $W=C_{2}\times C_{2}$
Case 1: $\Lambda_{0}$ is the $A_{1}\times A_{1}$ lattice, so that $\Lambda_{0}=\mathbb{Z}^{2}=\mathrm{Cub}$ is the standard lattice where the coordinates are $(x,y)$, and $W$ acts by reflection in the axes. We therefore find $\Lambda_{0}=\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$ where $\tilde{\mathbb{Z}}_{x}$ is the $x$-axis and $\tilde{\mathbb{Z}}_{y}$ is the $y$-axis. Factoring out the reflection in the $x$-axis or the $y$-axis gives two Serre spectral sequences. Comparing the two we see the differentials are zero in degrees affecting $H^{i}(W;\mathbb{Z}_{x})$ with $i\leq 3$ and hence $H^{2}(W;\tilde{\mathbb{Z}}_{x})\cong\mathbb{Z}/2$ and $H^{3}(W;\tilde{\mathbb{Z}}_{x})\cong(\mathbb{Z}/2)^{2}$. This gives the tabulated value. We obtain examples which are products of $SU(2)$ and $SO(3)$; the large number arise since there is a choice about which element of $W$ lifts to something whose square is non-trivial in each factor.
[…]
@@ 7.C.3. $W=C_{4}$
[…]
@@ 7.C.4. $W=D_{8}$
Considering the extension $1\longrightarrow\langle a,b\rangle\longrightarrow D_{8}\longrightarrow C_{2}\longrightarrow 1$ we see that $H^{i}(D_{8};\Lambda_{0})$ is $022X$ where $X$ is either of order 4 or 8.
Considering the extension $1\longrightarrow C_{4}\longrightarrow D_{8}\longrightarrow C_{2}\longrightarrow 1$, we see that $H^{i}(D_{8};\Lambda_{0})$ is $022X$ where $X$ is of order 2 or 4. Combining these we see that $X$ is of order 4.
@@ 7.C.5. $W$ of order 3
$0\longrightarrow\mathbb{Z}\longrightarrow\mathbb{Z}W\longrightarrow\mathbb{Z}[\omega]\longrightarrow 0.$
@@ 7.C.6. $W=D_{6}$
The one dimensional representations are again $\mathbb{Z}$ or $\tilde{\mathbb{Z}}$. If $\Lambda$ is of rank 2 then either $C_{3}$ acts trivially (and $\Lambda$ is inflated from the group $\Sigma_{3}/C_{3}$ of order 2) or else $M$ is a copy of $\mathbb{Z}[\omega]$ as a $C_{3}$-module. The reflection $\tau$ in the $x$-axis cannot act as a scalar (since $\tau g\tau=g^{-1}$), so it must have eigenvalues +1+1 and 1-1. The first example is when $\tau$ acts as complex conjugation on $\mathbb{Z}[\omega]$.
@@ 7.D. $W$ of prime order
We insert one more generic answer to hint at the complexities in higher rank. By [3, 74.3], the indecomposable representations of a group of prime order $p$ are $\mathbb{Z}$, $\mathfrak{a}$ for a fractional ideal $\mathfrak{a}$ in $\mathbb{Q}W$ (of rank p1p-1) or $\Lambda_{\mathfrak{a}}$ of rank $p$, where
@@
if $\Lambda_{\mathfrak{a}}=\mathfrak{a}\oplus\mathbb{Z}\cdot y$ additively, we may suppose $gy=1+y$. This means that $\mathbb{T}$ is a product of tori of dimension 1,p11,p-1 and $p$.
@@ 8. The case $W=C_{2}$, $\Lambda_{0}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$
This covers the groups $T\times O(2)$ and $T\times Pin(2)$.
@@ 8.B. Lattices and Weyl groups
(a) The lattice $\Lambda_{+}(m)$ with basis $\lambda=(m,0)$ for an integer $m\geq 1$. As a $W$-module this is $\mathbb{Z}$. It corresponds to $S=C_{m}\times\tilde{T}$ which has Weyl group the circle $Z/C_{m}$.
(b) The lattice $\Lambda_{-}(n)$ with basis $\lambda=(0,n)$ for an integer $n\geq 1$. As a $W$-module this is $\tilde{\mathbb{Z}}$. It corresponds to $S=Z\times C_{n}$ and has Weyl group $D_{4n}/D_{2n}$ of order 2.
Proof : The only comment to make is that if $\Lambda$ is an invariant rank 1 lattice then $g\Lambda=\Lambda$ so that a generator $\lambda$ is taken to $\pm\lambda$. For the Weyl groups we note $M(m,0)^{t}=(0,0)^{t}$ but $M(0,n)^{t}=(0,-2n)^{t}$. ∎
@@
Proof : Every rank 2 lattice $\Lambda$ contains an element on the $x$-axis, and we may choose the smallest positive element $(m,0)$, and similarly $(0,n)$ is the smallest positive point on the $y$-axis. If $\Lambda$ is generated by $(m,0)$ and $(0,n)$ we have a Type 1 lattice. Otherwise there is another element $(a,b)$ with $0<a<m,0<b<n$. Since g(a,b)=(a,b)g(a,b)=(a,-b) we find $(2a,0)$ is in the lattice so $2a=m$. If $2b$ is not a multiple of $n$ this contradicts the choice of $n$. Hence $(a,b)=(m,n)/2$ and we have a Type 2 lattice. ∎
@@ 8.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@
In the first case (corresponding to $\Lambda_{-}(n)$), $T^{2}/S$ is a circle with nontrival action so the cohomology of $T^{2}/S$ is in even degrees, and the lift of $\epsilon(G)$ is unique is unique up to conjugacy if it exists. In the split case $\epsilon(G)=0$ so the lift exists. In the non-split case, it lifts if and only if $n$ is even.
@@ 8.D. Summary
(this is the reason for indexing the lattice $\Lambda_{2}(2m,2n)$ using $(2m,2n)$, rather than $(m,n)$). Recalling $\Lambda_{2}(2m,2n)$ is generated by (m,n),(m,n)(m,n),(m,-n) one may check that the inclusions are both cofree and that the first is cofree and $W$-cotrivial. They are therefore represented vertically and horizontally in the picture.
@@
${H(\infty,1)}$${H(\infty,2)}$${H(\infty,3)}\textstyle{\cdots}$${H(\infty,\infty)}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}$${H_{1}^{s/ns}(4,1)}$${H_{1}^{s/ns}(4,2),H_{2}(4,2)}$${H_{1}^{s/ns}(4,3)}\textstyle{\cdots}$${H^{s/ns}(4,\infty)}$${H_{1}^{s,\pm}(3,1)}$${H_{1}^{s,\pm}(3,2)}$${H_{1}^{s,\pm}(3,3)}\textstyle{\cdots}$${H^{s,\pm}(3,\infty)}$${H_{1}^{s/ns}(2,1)}$${H_{1}^{s/ns}(2,2),H_{2}(2,2)}$${H_{1}^{s/ns}(2,3)}\textstyle{\cdots}$${H^{s/ns}(2,\infty)}$${H_{1}^{s,\pm}(1,1)}$${H_{1}^{s,\pm}(1,2)}$${H_{1}^{s,\pm}(1,3)}\textstyle{\cdots}$${H^{s,\pm}(1,\infty)}$
@@
${T\times D_{2}}$${T\times D_{4}}$${T\times D_{6}}\textstyle{\cdots}$${T\times O(2)}$${(C_{4}\times D_{2}),(C_{4}\times_{2}Q_{4})}$${(C_{4}\times D_{4}),(C_{4}\times_{2}Q_{8}),(C_{4}\times_{2}D_{4})}$${(C_{4}\times D_{6}),(C_{4}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{4}\times O(2)),(C_{4}\times_{2}Pin(2))}$${(C_{3}\times D_{2})^{2}}$${(C_{3}\times D_{4})^{2}}$${(C_{3}\times D_{6})^{2}}\textstyle{\cdots}$${(C_{3}\times O(2))^{2}}$${(C_{2}\times D_{2}),(C_{2}\times_{2}Q_{4})}$${(C_{2}\times D_{4}),(C_{2}\times_{2}Q_{8}),(C_{4}\times_{2}D_{4})}$${(C_{2}\times D_{6}),(C_{2}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{2}\times O(2)),(C_{2}\times_{2}Pin(2))}$${(C_{1}\times D_{2})^{2}}$${(C_{1}\times D_{4})^{2}}$${(C_{1}\times D_{6})^{2}}\textstyle{\cdots}$${(C_{1}\times O(2))^{2}}$
@@
8.E. Component structure
We read the component structure off Lemma 8.1 and Lemma 8.2. There is a trivial group over all subgroups on hte right hand vertical (those containing $\tilde{T}$) and all those corresponding to Type 2 lattices. Over all others it is a group of order 2.
@@ 9.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@ 10.A. Preamble
We have $W$ of order 2 and the toral lattice is $\Lambda_{0}=\mathbb{Z}W$. In this case $H^{2}(W;T^{2})=H^{3}(W;\Lambda_{0})=0$, and all such extensions split. To be concrete $W$ acts on $T^{2}$ by exchanging factors: this example is realised by the normalizer of the maximal torus in $U(2)$. The diagonal circle $Z$ is central and the antidiagonal $\tilde{T}$ consists of pairs $(z,z^{-1})$.
@@ 10.B. Lattices and Weyl groups
(a) The lattice $\Lambda_{+}(m)$ with basis $\lambda=(m,m)=m(1+w)$ for an integer $m\geq 1$. As a $W$-module, this is $\mathbb{Z}$. It corresponds to $S$ with $S_{e}=\tilde{T}$ and $S\cong C_{m}\times\tilde{T}$ with $\tilde{T}$. The Weyl group is and has Weyl group $Z/C_{m}$.
(b) The lattice $\Lambda_{-}(m)$ with basis $\lambda=(m,-m)=m(1-w)$ for an integer $y\geq 1$. As a $W$-module, this is $\tilde{\mathbb{Z}}$. This corresponds to $S$ with $S_{e}=Z$ and $S\cong Z\times C_{m}$, and has Weyl group of order 2.
@@
Type 2: If $m+n$ and mnm-n are even we have $\Lambda_{2}(m,n)$ which contains $\Lambda_{1}(m,n)$ as a sublattice of index 2. It has basis $\lambda_{1}=((m+n)/2,(m-n)/2)=((m+n)+(m-n)w)/2$ and $\lambda_{2}=w\lambda_{1}=((m-n)/2,(m+n)/2)=((m-n)+(m+n)w)/2$ and is $\mathbb{Z}W$ as a $W$-module. We have containments
@@
Proof : The eigenspaces of $\mathbb{Z}W$ are generated by $(1+w)$ and (1w)(1-w). If a lattice $\Lambda$ is generated by its intersections with $\mathbb{Z}W_{+}$ and $\mathbb{Z}W_{-}$, then it is of Type 1. Otherwise there will be an additional element, and we argue as in Lemma 8.2 to see that the lattice must be of Type 2. ∎